OpenStudy (anonymous):

A painter accidentally drops a can of paint from a high scaffold. How far has the can fallen after 4.0 s? a. -50 m c. -112 m b. -78 m d. -27 m

OpenStudy (rane):

is this all the information given ?

OpenStudy (anonymous):

yes, sadly. I am taking online courses. They really don't explain anything. I just take pretests and hope to pass them or i go through a pointless lesson watching videos of people who cannot explain very well.

OpenStudy (rane):

i don't know, there is no way this cn be solved with this info given

OpenStudy (anonymous):

And no information is given before pretests.

OpenStudy (anonymous):

Its all over Freefall.

OpenStudy (anonymous):

Most i can tell you.

OpenStudy (rane):

at least i can't all we have is t = 4s g= -9.8m/s^2

OpenStudy (anonymous):

I've got a freefall equation i found just now we may be able to use.

OpenStudy (anonymous):

V = sqrt(initial Velocity^2 + 2 x acceleration x distance)

OpenStudy (issy14):

Here you go: d=Vot+1/2at^2 that's the formula Your initial velocity is 0ms/ Starts at rest, your acceleration is 9.81m/ss. Plug in and you should get the answer. 78.4 round it up to 78

OpenStudy (fifciol):

it's very simple: $$h=\frac{gt^2}{2}$$

OpenStudy (rane):

did u drived that from some other equation ? @Fifciol

OpenStudy (anonymous):

Thank you @Issy14 and @Fificiol But I'll try both to see what i get.

OpenStudy (issy14):

Okis, you are welcome!

OpenStudy (anonymous):

@Issy14 Quick question, i'm new to those acronyms. Which ones plug in where and what are they called?

OpenStudy (issy14):

ok, Vo is initial velocity. so there you plug in 0m/s because the paint wasn't moving before it fell down, therefore it was at rest. d= 0m/s + 1/2 (9.81m/ss)(4)^2

OpenStudy (issy14):

Acceleration in this case is 9.81m/ss its going towards the earth and with gravity which is 9.81 therefore is positive.

OpenStudy (fifciol):

From the equation of motion with constant acceleration $$x=x_0+v_0t+\frac{gt^2}2{}$$ $$v_0=0 $$ and $$x_0=0$$ if you choose your origin clever enough

OpenStudy (anonymous):

@Issy14 Would i multiply 4 and 9.81 by ^2 ???? and then divide by 2?

OpenStudy (issy14):

no, the carat "^" means elevated, as in exponent. so is 4^2 = 16 then you multiply that by the acceleration which is 9.81m/ss after you multiply that you divide it by 1/2

OpenStudy (anonymous):

Yeah thats what i meant. And ahhh I See.

OpenStudy (anonymous):

I got it right with that equation. But hey @RANE Couple of your answers were incorrect XD

OpenStudy (issy14):

I'm glad!! Physics is a funny subject =).

OpenStudy (anonymous):

@Issy14 you're telling me! XD

OpenStudy (anonymous):

so what is it a b c or d

OpenStudy (anonymous):

wait what equation did you use? @therealjuicyj